# LUMINA-Net: Low-light Upgrade through Multi-stage Illumination and Noise Adaptation Network for Image Enhancement

### v1

@@ IV-A Experimental Datasets
IV-B Implementation Details
We implement LUMINA-Net using PyTorch. During training, images are randomly cropped to a size of $256\times 256$ to account for spatial variations. A batch size of 1 is applied for efficient memory usage, given the high-resolution image processing requirements. The ADAM optimizer [56] is employed with an initial learning rate of $1\times 10^{-4}$ , along with a cosine annealing learning rate scheduler for gradual decay. The number of training epochs is set to 400 to ensure convergence. For low-light enhancement scenarios, we empirically set the default correction factor to $\lambda=0.2$, as suggested in [32]. For the LOL dataset, the correction factor was adjusted to $\lambda=0.14$. The hyper-parameters $w_{0}$, $w_{1}$, $w_{2}$ and $w_{3}$ in the loss formulation are set to $w_{0}=5$, $w_{1}=1$, $w_{2}=1$ and $w_{3}=0.1$ based on empirical evaluation for optimal performance. Our models are run on NVIDIA TITAN XP GPUs.
IV-C Comparison with state-of-the-arts methods
@@ IV-D Quantitative Comparisons
Supervised methods utilize paired low-light and reference images to learn more precise mappings, yielding superior results. Despite being unsupervised, LUMINA-Net outperforms many supervised methods on both datasets. Its superior PSNR, SSIM, and LPIPS scores highlight its ability to preserve details, ensure color accuracy, and reduce artifacts, demonstrating its robustness in low-light enhancement without the need for reference images.
IV-E Visual comparisons
Figure 3 highlights the comparative performance of various low-light image enhancement methods on the LOL-real and SICE datasets. LUMINA-Net outperforms other methods by producing natural and visually balanced results with accurate brightness, color reproduction, and contrast. However, RetinexNet suffers from overexposure, which washes out image details and disrupts the natural appearance of colors, resulting in a loss of important features. This limitation is also reflected in RetinexNet’s metric values, as it achieves lower performance on metrics that emphasize detail preservation and structural similarity, such as SSIM and LPIPS.
When evaluating the similarity between the enhanced image and the ground truth, LPIPS (Learned Perceptual Image Patch Similarity) is the most indicative metric. Unlike PSNR and SSIM, which focus on pixel-level accuracy and structural similarity respectively, LPIPS is designed to measure perceptual similarity by comparing high-level features extracted from a deep neural network. This makes it more aligned with human perception, highlighting the issues in RetinexNet’s overexposed results. LUMINA-Net consistently achieves better LPIPS values, demonstrating its ability to produce outputs that closely resemble the ground truth in both appearance and perceptual quality.
Other methods, such as ZeroDCE and RUAS, fail to enhance extremely dark regions adequately, often leaving noise and artifacts. Similarly, KinD and URetinexNet, while effective in specific datasets, lack robustness in diverse lighting conditions. LUMINA-Net, in contrast, excels across all metrics, combining detail preservation, structural similarity, and perceptual quality to deliver realistic, high-quality results that outperform competing methods in both quantitative and qualitative evaluations.
IV-F Ablation Studies
To validate the effectiveness of LUMINA-Net’s components, including its modules and overall design, we conducted ablation experiments on the LOL dataset, with results presented in Table II and Figure 4.
In Figure 4 (a), removing the Over-Exposure Correction (OEC) Module caused overexposure and uneven brightness, significantly lowering PSNR and SSIM scores, emphasizing its importance in brightness management, (b) shows that excluding the Channel-Guided (CG) Module resulted in a loss of reflectance detail and structural consistency, underlining its role in reflectance-illumination separation for sharper details, (c), removing the Color Enhancement (CE) Module led to color distortions and reduced vibrancy, highlighting its importance in color accuracy. Finally, (d), our enhanced result, demonstrates the full LUMINA-Net configuration, showing preserved fine details, accurate colors, and balanced exposure. The results confirm that OEC, CG, and CE Modules are crucial for achieving high-quality low-light image enhancement, with reflectance enhancement being essential for maintaining sharpness and detail.
@@ IV-F Ablation Studies
w/o OEC 19.70 0.648
w/o CG 21.50 0.624
w/o CE 22.69 0.784
Ours 23.67 0.814
V Conclusion
In this paper, to address the challenges faced by existing low-light image enhancement methods, such as preserving image details in dark regions and accurately recovering colors, we propose LUMINA-Net. This method combines three key modules: the Channel-Guided (CG) Module for enhancing reflectance in dark regions, the Color Enhancement (CE) Module for adjusting illumination and ensuring accurate color restoration, and the Over-Exposure Correction (OEC) Module for handling overexposure. These modules work together to significantly improve both the quality and the naturalness of enhanced images. Through extensive experiments on low-light datasets, LUMINA-Net consistently demonstrates superior performance over existing state-of-the-art methods, achieving better detail preservation, accurate color recovery, and reduced artifacts, making it a promising solution for real-world low-light imaging applications.
Acknowledgment

### v2

@@ III-A Experimental Datasets
III-B Implementation Details
We implement LUMINA-Net using PyTorch. During training, images are randomly cropped to $256\times 256$, and a batch size of 1 is used for efficient memory usage. The ADAM optimizer [44] is employed with an initial learning rate of $1\times 10^{-4}$, along with a cosine annealing scheduler. Training runs for 400 epochs for convergence. The default correction factor is $\lambda=0.2$, as in [37], and adjusted to $\lambda=0.10$ for the LOL dataset. The loss weights $w_{0}=5$, $w_{1}=1$, $w_{2}=1$, and $w_{3}=0.1$ are set based on empirical evaluation. Models are run on NVIDIA TITAN XP GPUs.
III-C Comparison with state-of-the-arts methods
@@ III-D Quantitative Comparisons
III-E Visual comparisons
Fig. 3 compares various low-light image enhancement methods on the LOL-real and SICE datasets. LUMINA-Net outperforms other methods by producing natural, balanced results with accurate brightness, color, and contrast. In contrast, RetinexNet suffers from overexposure, washing out details and disrupting color reproduction, leading to lower performance in metrics like SSIM and LPIPS.
LPIPS (Learned Perceptual Image Patch Similarity) is a key metric that captures perceptual similarity using deep neural network features. Unlike PSNR and SSIM, LPIPS better aligns with human perception, exposing overexposure issues in methods like RetinexNet. LUMINA-Net achieves lower LPIPS scores, reflecting closer visual similarity to the ground truth.
In contrast, methods like ZeroDCE and RUAS struggle with dark regions and noise, while KinD and URetinexNet lack robustness under varied lighting. LUMINA-Net outperforms across all metrics by preserving details, enhancing structure, and delivering high perceptual quality.
III-F Ablation Studies
To validate the effectiveness of LUMINA-Net’s components, including its modules and overall design, we conducted ablation experiments on the LOL dataset, with results presented in Table II and Fig. 4.
@@ III-F Ablation Studies
w/o OEC 19.70 0.646
w/o CG 21.50 0.622
w/o CE 22.69 0.782
Ours 23.92 0.812
In Fig. 4 (a), removing the OEC Module caused overexposure and uneven brightness, significantly lowering PSNR and SSIM scores, emphasizing its importance in brightness management, (b) shows that excluding the CG Module resulted in a loss of reflectance detail and structural consistency, underlining its role in reflectance-illumination separation for sharper details, (c), removing the CE Module led to color distortions and reduced vibrancy, highlighting its importance in color accuracy. Finally, (d), our enhanced result, demonstrates the full LUMINA-Net configuration, showing preserved fine details, accurate colors, and balanced exposure. The results confirm that OEC, CG, and CE Modules are crucial for achieving high-quality low-light image enhancement, with reflectance enhancement being essential for maintaining sharpness and detail.
IV Conclusion
In this paper, we introduce LUMINA-Net to address challenges in low-light image enhancement, such as preserving details in dark regions and accurately recovering colors. It integrates three modules: CG for reflectance enhancement, CE for color restoration, and OEC for overexposure correction. LUMINA-Net improves image quality and naturalness, outperforming existing methods in detail preservation, color accuracy, and artifact reduction.